OpenStudy (anonymous):

MEDAL WILL BE AWARDED - I JUST NEED YOUR ANSWER PLEASE to see if it matches mine - For a curve with equation y = 3x2 − x: find the gradient of chord PQ, where P is the point (1, 2) and Q is the point ((1+h),3(1+h)2 −(1+h))

OpenStudy (wolf1728):

In order to get the gradient or slope we will need the exact position of Q. What is 'h' ?

OpenStudy (anonymous):

arrow 0

OpenStudy (anonymous):

h->0 i would assume because it doent state @wolf1728

OpenStudy (wolf1728):

Okay I made a graphic with point P labelled but again where is point Q?

OpenStudy (anonymous):

what s your answer ?

OpenStudy (anonymous):

@Myininaya

OpenStudy (anonymous):

@Satellite

OpenStudy (anonymous):

i just need to check my answer does anyone have an answer :(

OpenStudy (campbell_st):

1st question, what answer do you have...?

OpenStudy (anonymous):

its only one question but i got 3h+1 and the answer is meant to be 3h + 5

OpenStudy (campbell_st):

well the answer is correct... I think the problem is the way you have distributed \[\frac{2 - [3(1+h)^2 -(1 + h)]}{1 - (1 + h)} = \frac{2 - (3 + 6h + 3h^2 - 1 - h)}{h}\] hope this helps

OpenStudy (anonymous):

ill try it thanks @cambell_st

OpenStudy (anonymous):

now i get -6h-9 what did you get?

Directrix (directrix):

I got 3h + 5 which @campbell_st said is correct. Take note of what he said about the way you used the distributive property.

OpenStudy (anonymous):

could you possibly show me your full working

OpenStudy (campbell_st):

I missed the negative in the denominator above, a typo so you get \[\frac{2-(2 + 5h + 3h^2)}{-h} = \frac{2 -2 -5h -3h^2}{-h} = 5 + 3h\]